Displaying reports 681-700 of 1089.Go to page Start 31 32 33 34 35 36 37 38 39 End
Reports until 12:35, Wednesday 29 May 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 12:35, Wednesday 29 May 2024 - last comment - 16:01, Wednesday 29 May 2024(11616)
AOM debug

[Lee, Torrey]

-Low voltage high amp draw from RF power source

-Checking the power of the RF drive on the aom. Within limits with some loss through BNC and factor of two from termination.

-Tested everything. The homeade power source was shorted.

-Re-soldered power source, plugged everything in, and we have frequency shifted light. More to follow after lunch.

Images attached to this report
Comments related to this report
Torrey Cullen - 16:01, Wednesday 29 May 2024 (11617)

-Installed 1/4 waveplate and mirror to reflect back through AOM.

-Align to see the twice shifted beam. Can confirm this beam is from the difracted beam as both dissapear if you turn off the AOM RF drive. Also can use iris to block the normal beam and pass the first order diffracted beam.

-There is VERY little clearance for the beam due to the AOM RF drive cable. The cable is almost exactly at 4 inches from the breadboard. Thinking about using an AOI greater than 45 degrees to align into fiber. The thorlabs BB1-EO3 shows good reflectance at low AOIs but not at large ones.

-The current configuration on the table can be seen here. The twice shifted beam is fairly astigmatic so alignment can most likely be improved.

Images attached to this comment
GQuEST Output Filter Cavities
Torrey Cullen - posted 10:31, Wednesday 29 May 2024 (11614)
Preparing for coresonant cavity

Because the location of the curved super optic has changed inside the cavity from its initial position, we must change the mode matching for the 1550nm light in the filter cavity. Previously, the waist was calculated to be at the third mirror the 1550nm light encounters, whereas now the waist is on the input coupler. The new MM solution is approximately f1 = .250 @ .902m and f2 = 1.5 @ 1.546m, where the distances are counting from the 1550nm fiber collimator. I placed new lenses according to the solution and measured the beam to be slightly too big at the measure point. This means L1 was too far back, it was moved forward <1 inch to match the predicted beam size. We can see the first pass beam on the SWIR camera but alignment will need to be improved before we see any coresonance.

Images attached to this report
GQuEST SHG
Torrey Cullen - posted 13:05, Friday 24 May 2024 (11613)
AOM initial power on

[Torrey, Ian]

-Noticeable bump on SHG sled, things were realigned.

-Decent initial alignment into AOM. Going for power on now. 24 V, 0.9 A DC supply per the requirements of the 2 W amplifier we have. https://www.minicircuits.com/pdfs/ZHL-1-2W+.pdf 

-According to RF amplifier the absolute input Vpp is 2 V (10 dBm). We are going to start with a 20th of this, 0.1Vpp. This should correspond -16 dbm to start with a ~35 dbm boost giving 19 dbm. This corrersponds to 0.080W.  No observed diffraction.

-Bump it up to 0.317 Vpp. This is -6 dbm. +33 dbm boost min gives 27 dbm, which is .5 W. No observed diffraction. Going to change alignment now.

-No visible difference when changing alignment on PY005. Need a different strategy for now. Not sure if we're under driving the AOM either. RF input signal is powered off and disconnected for now.

Controls FPGA
Jeffrey Wack - posted 09:45, Friday 24 May 2024 - last comment - 11:58, Friday 31 May 2024(11611)
Connecting to Red Pitaya with PYNQ image

First, I downloaded the image found here (DSP sandbox PYNQ-Red Pitaya) and wrote it to a micro SD card. After that I connected the Red Pitaya to:

1. Power (micro USB)

2. Serial connection to computer (micro USB)

3. Ethernet connection to router

Then, I went to "http://pynq:9090" I was met with a jupyter notebook password screen. The password is "xilinx".

In the jupyter menu you can open a terminal to access the red pitayas Linux system. I followed the instructions here (PYNQ getting started) to change the host name to "pynqpitaya"

I ran the command "ifconfig". There are multiple IP addresses. Why?

The address that allows one to ssh or scp onto this red pitaya is 192.168.50.14, which corresponds to "eth0" rather than "eth0:1"

After using scp to get the analog echo bitfile and notebook from Chris' logpost onto the board, I ran the notebook. The first cell runs successfully, and a blue light comes on the red pitaya. However the cell with the write statement never executes. When I measure the DAC outputs on an oscilloscope, they measure about +1V and -1V. No response is seen to an input pulse on ADC 1.

 

Comments related to this report
Jeffrey Wack - 11:58, Friday 31 May 2024 (11619)

Today I tried this procedure again with a new SD card, and the result was the same. There is still no echo from the Red Pitaya. Here are three possible explanations:

1) The Red Pitaya I have is broken. I don't like this explaination because I am able to log into the Jupyter notebook and run the first cell without any issue.

2) There is something wrong with my procedure for writing the PYNQ image to an SD card which I have repeated each time. Like the last idea, this seems implausible due to the fact that the booting the system and accessing the jupyter notebooks works fine.

3) there is some difference, either between the bit file and the board, or between the bit file and the PYNQ image, which is causing things to work differently than on Chris' system. My board is a Red Pitaya STEMlab 125-14 LN. The PYNQ image is Pynq-Redpitaya-125-14-3.0.1.img

GQuEST SHG
Torrey Cullen - posted 16:49, Thursday 23 May 2024 (11607)
AOM progress

[Torrey, Daniel, Sander]

Continuing AOM progress. A f= 100mm lens is in place, with the AOM, and power supply ready. We ran into the problem of the input for the AOM being SMB, of which we do not have an adapter for. This will be on hold until we can procure one from another lab. Digikey sells these but have a long lead time. Amazon also has these.

Equipment General
Torrey Cullen - posted 16:47, Thursday 23 May 2024 (11606)
3rd Basler camera now in use

[Torrey, Sander, Daniel]

Third basler camera is now in use and is on the SHG sled (SN 24839320). Tested as is working.

Equipment General
Torrey Cullen - posted 13:41, Thursday 23 May 2024 (11605)
RF Power Amplifiers

Current list of RF power amplifiers we have in B102:

-Mini circuit 5 W amp x2

-Mini Circuits 2 W amp x2 (one in use on SHG sled)

-ENI 3W amplifier x1 

These are all on the top shelf of the beige cabinet to the left of the toolbox.

Images attached to this report
Controls Optimal Controls
Ian MacMillan - posted 13:17, Thursday 23 May 2024 (11604)
Improved RMS from Buzz control for DHARD Yaw

Attached is the plot for a number of controllers calculated by buzz for the ASC DHARD Yaw DOF. The plot's axis represent the total RMS on the Y-axis in Radians and a weighted RMS on the X-axis that is proportional to the BNS detection range. This is just an estimate using the square of the BNS FOM that I was using previously. The number by each point represents the phase margin of each controller. The outline color is also supposed to represent this but there is some bug that gets the colors slightly wrong from the scale. I will get to fixing this. A recent hand tuned controller is also present for comparison. The plot was calibrated using the 5.2e-11 rad/ct from Elenna's Alog post on ASC Calibration. The hand tuned controller shows an RMS of approximately 1.6 nrad RMS. given that our model represents is modeling a day with higher than normal low frequency noise this compares favorably with the 0.36 nrad RMS measured recorded in the same post by Elenna. This indicates that our model is a good representation of the system, at least in this limited way. Given the calibration it is possible to get a controller that with similar phase margin with an order of magnitude less overall RMS.

The next step is to add the updated BNS FOM so that the BNS RMS value has a physical representation. Currently it seems that the updated BNS FOM causes problems with the current solver.

Non-image files attached to this report
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 00:23, Thursday 23 May 2024 (11603)
Modified AOM Mount to mount the Gooch & Housego 3080-125 AOM on a Thorlabs PY005 5 axis Stage

To one of the AOM mounts we ordered from 3D Hubs, I added two tapped #6 holes to the following two locations, assuming the center of the mount is the origin: (+/- 0.875, 0.25). The y coordinate is not zero because there are existing holes at (+/- 0.75, 0) which would interfere.

 

The Gooch & Housego 3080-125 AOM fits very well onto these holes and it mounted in B102.

GQuEST SHG
Torrey Cullen - posted 16:57, Wednesday 22 May 2024 (11602)
Start of AOM work on SHG sled

[Torrey, Ian]

We have an AOM to use while we obtain our own. We have decided on using this to power it. Elie Bataille quotes it at 5W max so we will under drive this amplifier. We took an intial profile of what the beam looks like near where the AOM will be. Space is very limited to build a proper double lens solution (i.e. sum of the focal lengths = the distance between them) as seen in this. Because of this we are proposing a single lens solution where the mirror is placed approxiamtely at the focal length, seen here.

Images attached to this report
Equipment Loans (GQuEST)
Daniel Grass - posted 13:42, Wednesday 22 May 2024 - last comment - 11:55, Wednesday 30 October 2024(11601)
Gooch & Housego 3080-125 AOM Borrowed from the Endres Lab

I borrowed a Gooch & Housego 3080-125 AOM from the Endres Lab. My point of contact is Elie Bataille. It is currently in East Bridge B102.

Comments related to this report
Daniel Grass - 17:30, Thursday 23 May 2024 (11608)

I have also borrowed an SMB female to SMA and SMA to BNC male

Daniel Grass - 11:55, Wednesday 30 October 2024 (11950)

I returned the AOM and cable to Hannah Manetsch.

Controls FPGA
Jeffrey Wack - posted 12:33, Wednesday 22 May 2024 (11600)
Red Pitaya Attempted Quickstart

I followed the instructions here to flash this OS to a Red Pitaya and connect over ethernet. After inserting the SD card, connecting the ethernet cable and power, the Red Pitaya enters a reboot cycle, as indicated by the periodic turning off and on of the blue LED and the following red and orange flashes. The steady green light tells us there is no power issue. This info about the LED indicators was found on this FAQ, which suggests the rebbot cycle is due to a missing external clock signal.

GQuEST Output Filter Cavities
Torrey Cullen - posted 11:02, Wednesday 22 May 2024 (11599)
Updated OFC sled layout

Updated layout of the first output filter cavity. Note the new convention on labeling the actual filter cavity super optics mirrors. Serial numbers and mirror discription can be found at https://wiki.mccullerlab.com/DCC/S2400001 .

Images attached to this report
Non-image files attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 11:55, Tuesday 21 May 2024 (11596)
Robust Lock on OFC using 775 nm light

[Ian, Torrey]

Procedure notes in 775 cavity progress in the lab today:

-Improved alignment in vertical direction. Alignment is sufficient for initial tests and cavity characterization but can be improved. Mode scan suggests that mode matching could be improved. 

-Polarization fluctuations in fiber due to fans is problematic because of how we are attenuating the power (PBS). We have reduced the overall laser power now that alignment is done until 1811 PD in not saturated. This is done by turning the amplifier all the way down and removing the PBS. The 1811 is still saturated in this configuration. Temporarily changed the SHG polarization so less power goes into fiber on the SHG sled to the OFC sled.

-This fixed any problems with the refl PD we now have a robust lock (with the laser).

-Peak to peak error amplitude with no tank circuit and 5 mV set point offset is between 10.5 mV to 11 mV. Tank circuit was stolen from the 1550nm EOM as the capacitance of the 775nm EOM is the same (14 pF). (Note that we have the parts to build another LC circuit on the 1550nm path but are waiting for the little circuit boxes that should arrive with the big thorlabs order that has been placed.) Error amp is now 34 mV requiring an error set point of 11 mV. We did not adjust the modulation frequency of 50.68 mHz, could potentially get some more amplitude from this.

-The cavity was then locked with the laser and we took measurements for G and the noise spectrum at the error point in the same way as previous posts. Data for this can be found in the nextcloud: "\Nextcloud\GQuEST\B102\Output Filter Cavity\775nm locks\Initial laser lock"

 

Images attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 11:14, Tuesday 21 May 2024 - last comment - 11:24, Tuesday 21 May 2024(11595)
775 nm light output cavity lock

[Ian, Torrey]

We have aligned the 775nm light path and cavity mirrors to the point that we have a large a mount of 0,0 mode flashing! We did this by first removing the input coupler (OFCM3 see labeled photo in post 11575) from the cavity so we could see the beam inside the cavity without strong reflections. The using the 3D printed jigs to get the beam close to the proper alignment for the other three mirrors. After we had the alignment correct on those we installed the input coupler (OFCM3). Then we walked the beam around by moving the 775 nm input coupling mirror (OFCM3) and the one diagonal from that (OFCM2) until we saw two beams on the transmitted photodiode. Then we continued to walk it until we saw three. making these overlap on the photodiode made use see HOMs. We played with the alignment walking the beam until we saw the lowest order 0,0 mode flashing with the others.

After we saw the first mode flashing we installed a 50/50 beamsplitter on the transmitted path allowing us to use the Bassler camera that we had been using for alignment and installed a transmission photodiode (Thorlabs PDA50B2). This photodiode is only rated to go down to 800 nm but it has enough sensitivity for the basic lock. We will need to replace it when we get new PDs. We also realigned the reflected beam path because that was now off from our alignment. 

With all of those fixes and installations of new BNC cables (which we had to make out of a number of shorter ones. we should buy more) we were able to get a rough lock on the 0,0 mode shown in first attached image. We think that there needs to be adjustments to the mode matching to make the lock robust.

Images attached to this report
Comments related to this report
Torrey Cullen - 11:24, Tuesday 21 May 2024 (11597)

Post updated from earlier by Ian

Controls General (GQuEST)
Jeffrey Wack - posted 11:09, Friday 17 May 2024 - last comment - 10:53, Wednesday 29 May 2024(11582)
Block diagram for Pound-Drever-Hall stabilization

Behavior:

Attached is a block diagram describing the behavior of a DSP system permforming feedback control for the Pound-Drever-Hall (PDH) frequency stabilization technique. This block diagram is based on Liquid Instrument's laser lock box, which we have been using to lock our current cavity.

There are two modes of operation: 'scanning mode' and 'locking mode.' When the toggle (on the right side of the diagram) is down, we are in 'locking mode' and the control loop is closed.

In locking mode there are two main tasks. First, we must obtain the error signal from the incoming photodiode signal by demodulation. We mix the incoming signal with the modulation tone, (with an adjustable phase offset) then we low-pass filter to remove unwanted high-frequency signal coming from the mixing. Then we subtract the desired setpoint from this signal to obtain the error signal. At this location, the error point, we want to be able to take time series of the values. These timeseries will be used to evaluate the characteristics of the loop and to make the transition from scanning mode to locking mode. Next, the error signal is passed through a series of biquad filters (the controller) to create the signal to be sent to the actuators. A constant offset is added to this signal and it is sent out to the actuator.

In scanning mode there is no feedback. Instead, the actuator voltage is swept by the sawtooth generator, and the error signal is monitored at the error point. Using the timeseries of the error signal during this sweep we can adjust the value of the constant offset and toggle the switch, exiting sweep mode and closing the loop.

Specifications:

To meet the requirements of the GQuEST experiement, we want to lock the cavities with a 10kHz unity gain frequency, and a 'low enough' root-mean-squared displacement. We will continue to work to understand what is 'low enough' and how this translates to FPGA specs, but until then here is a list of our best guesses at the approximate desired stats. These are based on the stats of the Moku and a few calculations, some of which are found in this log post and its comments. In particular, as the filter sample rate is increased, the number of bits required for the filter coefficients will increase as well. 

Filter sample rate - 25MHz (Moku IIR filters use 40MHz)

ADC/DAC bit count - 16bits (Moku combines a 10 MSa/s, 18-bit ADC and a 5 GSa/s, 10-bit ADC)

Biquad internal bit count - 32bits (Moku uses 48bit filter coefficients)

Total input-output latency - 1us (Moku has around 1us)

Images attached to this report
Comments related to this report
Ian MacMillan - 11:18, Friday 17 May 2024 (11588)

Can you add a sampling point before and after the SOS so that I can analize how well they are preforming?

Chris Stoughton - 11:29, Friday 17 May 2024 (11589)

Jeff, this looks good.  What about timing stability?  How stable do frequencies need to be?  For more than one PDH control loop, are the feedback loops independent or do we need them to be synchronized?

 

Jeffrey Wack - 11:49, Friday 24 May 2024 (11612)

Attatched is a good reference on Pound-Drever-Hall stabilization

Non-image files attached to this comment
Jeffrey Wack - 10:53, Wednesday 29 May 2024 (11615)

Attached is a powerpoint by Matt Evans on digital filter implementations, including a low noise second order section.

Non-image files attached to this comment
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 20:29, Thursday 16 May 2024 - last comment - 13:58, Tuesday 04 June 2024(11586)
3D Printed a Holder for the Basler ace GigE C camera

I 3D printed a Holder for the Basler ace GigE C camera, which we are using to image the 775 nm light. The holder interfaces with 3 M3 screws on the bottom of the camera and allows for a #8 socket head cap head screw to be added. The bottom of the holder is 1 inch from the center of the camera. The holder has some slots to allow for air to flow below the camera for ventilation and hopefully keep it cooler. I'm not sure of the efficacy of this since the other 5 faces were open on the other camera and it still got quite hot (although still within spec).

Attached are SolidWorks and STL files for the part I made and the camera, plus a PDF design of the camera.

Non-image files attached to this report
Comments related to this report
Daniel Grass - 20:17, Thursday 23 May 2024 (11609)

I printed another mount with wider M3 holes for easier insertion. See attached files.

For a future version, perhaps the M3 holes shouldn't be counterbored so there is less filament to clean up. This would require the correct M3 screw length, which might be harder to find for a longer screw.

Non-image files attached to this comment
Daniel Grass - 13:58, Tuesday 04 June 2024 (11628)

I redesigned and made a mount that allows for a C mount to be attached to the front so that ND filters and lenses can be easily added.

See attached files.

Images attached to this comment
Non-image files attached to this comment
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 11:56, Tuesday 07 May 2024 - last comment - 20:20, Thursday 23 May 2024(11557)
Piezo Alignment Tool

[Torrey, Daniel]

Torrey thinks that the central axes of the piezo and the mirror are misaligned. I designed a part that aligns the small piezo and a 1/4 in thick spacer with a #8 through hole with the "piezo top". This piezo top should be well aligned with the piezo bottom that holds the mirror. I believe this should give alignment of the axes to within ~5 thou rms (3 thou from the piezo top to the base, 4 thou from the mirror in the SM1 threads, and ~1 thou from this tool).

Attached is the part file with CAM as well. I decided to make this part in a CNC Lathe for its precision compared to a 3D print. An important consideration is the radius of curvature of the cutting tool. This is why there is a notch toward the thickest part of the tool and why the levels of the tool don't match the levels of each part. If one were to 3D print this part, they should remove the notch so that there are no overhung sections.

The blue colored photo is the CAM simulation.

Images attached to this report
Non-image files attached to this report
Comments related to this report
Daniel Grass - 14:24, Tuesday 14 May 2024 (11571)

I also made an alignment tool for a larger Thorlabs ring piezo. This requires a larger spacer, so I made a 1/4 thick (actually ~0.24 in) ring with a 1 in OD and ~0.358 in ID, slightly larger than the 9 mm ID of the ring piezo. This ring is aluminum because I didn't want to machine steel because it's tougher. 

Attached is this file (medium) and STL files for this part and the other part.

I made this part on the CNC lathe and it fits well.

I also manually made a part to align the noliac piezo. This is essentially a 0.7" long rod with a 0.47" diameter and a 0.25" thick, 0.5" diameter cap so the rod doesn't go all the way into the piezo assembly. There isn't a part file for this.

This Noliac part was too large. The noliac ID appears to be closer to 0.463"

Non-image files attached to this comment
Daniel Grass - 20:17, Thursday 16 May 2024 (11584)

I made a piezo alignment tool for the Noliac NAC2125-H08. The diameter of this tool is 0.460" and it fits pretty smoothly around the piezo. Attached is the SolidWorks File (with CAM as well) and the file as an STL.

Non-image files attached to this comment
Daniel Grass - 20:20, Thursday 23 May 2024 (11610)

The Noliac piezo is a bit loose around the 0.46" diameter part, so I printed a conical version. I would have printed a version with a smaller conical angle, but I was worried about getting the part stuck because a 1-3 degree taper is prone to getting stuck.

Non-image files attached to this comment
Displaying reports 681-700 of 1089.Go to page Start 31 32 33 34 35 36 37 38 39 End